Common stock : The firm's common stock is selling at RM 60. The floatation cost is 3 percent of selling price. At present, the company's growth rate is 5 percent and the latest dividend paid was RM 0.90. Calculate the cost of each alternative if the company is in a 40 percent tax bracket

1
Expert's answer
2021-11-10T10:08:19-0500

D1 = is the dividend will be paid next year


D1 = DO "\times" (1+g)


DO = Dividend just paid, latest or latest dividend paid, last year dividend paid


D1 = 0.90 ( 1 + 0.05)

D1 = 0.90 ( 1.05)

D1 = 0.945 


D1 = 0.945

Selling price, SP = 60

Floatation cost, FC = 3"\%\times" 60 = 1.80

Growth rate, G = 0.05 OR 5"\%"


Effective rate="\frac{D1}{(SP – FC)} + G"


= "\frac{0.945}{(60– 1080)} + 0.05"

= "\frac{0.945}{58.2} + 0.05 = 0.06623\times100\% = 6.62\%"


The best alternative for the company is using Common Stock Financing because effective rate (6.62%) is lower.